The sum of Beth and Ashley's ages is 84. John's age is 12. The sum of all three ages is 120. Find their ages.
Answer by TimothyLamb(4379)

x = beth age
y = john age
z = ashley age
---
y + z = 84
y = 12
x + y + z = 120
---
put the system of linear equations into standard form
---
y + z = 84
y = 12
x + y + z = 120
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = beth age = 36
y = john age = 12
z = ashley age = 72
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---